# Towards the Training of Deeper Predictive Coding Neural Networks

### v2

@@ Abstract
Predictive coding networks trained with equilibrium propagation are neural models that perform inference through an iterative energy minimization process. Previous studies have demonstrated their effectiveness in shallow architectures, but show significant performance degradation when depth exceeds five to seven layers. In this work, we show that the reason behind this degradation is due to exponentially imbalanced errors between layers during weight updates, and predictions from the previous layer not being effective in guiding updates in deeper layers. We address the first issue by introducing two novel methods to optimize the latent variables that use precision-weighting to re-balance the distribution of energy among layers during the ‘relaxation phase’, and the second issue by proposing a novel weight update mechanism that reduces error accumulation in deeper layers. Empirically, we test our methods on a large number of image classification tasks, resulting in large improvements in test accuracy across networks with more than seven layers, with performances comparable to those of backprop on similar models. These findings suggest that a better understanding of the relaxation phase is important to train models using equilibrium propagation at scale, and open new possibilities for their application in complex tasks.
@@ 1 Introduction
Training large-scale models on GPUs is extremely expensive in terms of energy consumption. To address this problem, a recent direction of research is studying the use of alternative accelerators that leverage the properties of physical systems to perform computations (Wright et al., 2022; Momeni et al., 2024). A popular example is hardware that performs in-memory computations using memristor crossbars (Tsai et al., 2018; Haensch et al., 2018). However, transitioning to new hardware without altering the main algorithm — error backpropagation (Rumelhart et al., 1986) — has proven to be challenging due to two central issues: the requirement of sequential forward and backward passes, and the need to analytically compute gradients of a global cost function. These requirements necessitate digital hardware architectures that precisely match forward computations with their corresponding gradient calculations in low-noise environments, as even minor fluctuations can propagate numerical errors that can alter the final performance of the model.
To address the above problems, there is a direction of research that has studied learning algorithms for deep neural networks that rely on local computations only (Bengio, 2014; Hinton, 2022). One way of doing this is by training the model using equilibrium propagation, a framework that allows the learning of the parameters of a neural network by simulating a physical system brought to an equilibrium (Scellier and Bengio, 2017). This physical system is usually defined via an energy function that describes the state of a neural network in terms of its weights and neurons, with different functions describing different systems (Hopfield, 1982; Krotov and Hopfield, 2016).
@@
In recent years, researchers have devoted significant effort to scaling up the deployment of energy-based models. Two recent works have also carefully benchmarked multiple variations of the usual learning algorithms by using the Hopfield energy function (Scellier et al., 2024), and the predictive coding energy (Pinchetti et al., 2024), showing that this class of models is able to perform as well as standard deep learning models trained with backpropagation when it comes to training shallow models, with a maximum of five or seven layers. This result, however, does not translate to the case of deeper models, where we observe a significant degradation of performance. As most of the successes of modern deep learning rely on very deep architectures, understanding and addressing the causes of this degradation is then of vital importance if we aim to train large-scale predictive coding networks.
@@
It has recently been shown that in a three-layer model, the energy concentrated in a layer can be up to an order of magnitude larger than the energy concentrated in the layer before (Pinchetti et al., 2024). While such shallow models can still achieve good test accuracies, we conjecture that this ’energy imbalance’ becomes a critical bottleneck in deeper architectures, leading to performance degradation in a way that is conceptually similar to the vanishing gradient problem (Hochreiter, 1998). More precisely, this imbalance prevents the effective propagation of energy — and crucially, the associated error information — from the output layer back to the early layers, creating two problems: first, it prevents the model from fully leveraging its depth, as early layers receive insufficient error signals for effective training; second, latent states may diverge substantially from their forward pass values, due to the excessive energy in later layers, thereby resulting in sub-optimal weight updates.
Addressing this energy imbalance requires mechanisms that can adaptively modulate the energy of different layers as it propagates through the network — a challenge remarkably similar to what biological neural systems solve through precision-based regulation (Clark, 2013; Mumford, 1992). In the brain, the flow of information is in fact believed to be heavily influenced by precision weighting, a mechanism that dynamically rescales error terms, effectively balancing top-down and bottom-up signals between different brain regions (Rao and Ballard, 1999; Friston, 2005; Friston and Kiebel, 2009; Clark, 2013). Despite its central role in biological systems, precision weighting has been largely overlooked in machine learning applications, and often set to $1$ in experimental setups when considered in algorithmic descriptions (Whittington and Bogacz, 2017; Salvatori et al., 2023; Song et al., 2020). Inspired by this neuroscientific principle, we propose to leverage precisions to regularize energy propagation in predictive coding networks, aiming to improve their performance in image classification tasks by addressing the identified energy imbalance. We do so by first analyzing the energy propagation of deep convolutional architectures, and then propose time-dependent precisions that address them. The results show that this largely improves test accuracies, hence confirming our conjecture of a causal relation between energy propagation and empirical results. Our contributions are briefly as follows:
  • We empirically study the propagation of the energy on different layers of the model, and show that such energy is orders of magnitudes larger in the layers closer to the label, and almost non existent in the first layers. This strongly supports the hypothesis that the energy of deep models fail to reach the first layers. We also show that a similar phenomenon is not observed in high-performing models trained with backpropagation.
  • To address this imbalance, and test on whether it impacts the final test accuracy of the model, we propose two different combinations of precision-weighting, which depend on both time and layer number. The first proposal, called spiking precision leverages very large precisions as soon as the energy reaches a specific layer, to boost it towards the following layer. The second proposal, which we call decaying precision, introduces an exponential decay in precisions that heavily penalize later layers. The results show that both methods regulate the energy imbalance, with the spiking precision providing the largest improvements in terms of test accuracies in deep models. In addition, we also propose a variation of batch normalization (Ioffe and Szegedy, 2015), which further improves the results under the novel kinds of inference.
  • To further boost the results, we introduce a novel weight update mechanism that changes the way weight parameters are updated. This method uses both the predictions computed at initializations (hence adding a degree of implausibility), and the neural activities at convergence, to compute the updates. Again, the results show that this largely improves the performance of predictive coding networks in models with more than ten layers. Furthermore, combining our novel weight updates and precision-weighed inference allows us to reach competitive results against backprop on challenging benchmarks for the biologically-plausible community, such as the training of convolutional models with $15$ layers on Tiny ImageNet.
@@ 2 Related Works
EP is a learning algorithm for supervised learning that is largely inspired by contrastive learning on continuous Hopfield networks (Movellan, 1991). Here, neural activities are updated in two phases: In the first, to minimize an energy function defined on the parameters of the neural network; in the second, to minimize the same energy with the addition of a loss function defined on the labels (Scellier and Bengio, 2017). Interestingly, these two phases allow us to approximate the gradient of the loss function up to arbitrary levels of accuracy using finite difference coefficients (Zucchet and Sacramento, 2022). The consequence is that EP can be seen as a technique that allows the minimization of loss functions using arbitrary physical systems that can be brought to an equilibrium, and it has hence been studied in a large number of domains (Scellier, 2024; Kendall et al., 2020). In terms of simulations that aim to scale up machine learning experiments, most of the works performed experiments using Hopfield energies (Hopfield, 1982), mostly on image classification tasks using convolutional networks (Laborieux et al., 2021; Laborieux and Zenke, 2022). The state of the art is that EP models are able to match the performance of BPTT (BP Through Time) on models with $5$ and $7$ hidden layers (Scellier et al., 2024), with the exception of hybrid models, which manage to achieve a good performance on models with $15$ layers by alternating blocks of layers trained with BP and blocks trained with EP (Nest and Ernoult, 2024).
@@
The formulation of PC that we use here was initially developed to model hierarchical information processing in the brain (Rao and Ballard, 1999; Friston, 2005). Intuitively, this theory states that neurons and synapses at one level of the hierarchy are updated to better predict the activities of the neurons of the layers below, and hence minimize the prediction error. Interestingly, the same algorithm can be used as a training algorithm for deep neural networks (Whittington and Bogacz, 2017), where several similarities with backpropagation were observed (Song et al., 2020; Salvatori et al., 2022a). To this end, it has been used in a large number of machine learning tasks, from image generation and classification to natural language processing and associative memory (Sennesh et al., 2024; Salvatori et al., 2023; Pinchetti et al., 2022; Ororbia and Kifer, 2020; Salvatori et al., 2021). Again, the state of the art has been reached by training convolutional models with $5$ hidden layers, with performance starting to get worse as soon as we use models that are $7$ layers deep (Pinchetti et al., 2024). The connection between PC and EP is well explained by the concept of bi-level optimization (Zucchet and Sacramento, 2022), where the neural activities used for learning are the equilibrium state of a physical system. For a description of a more general framework that describes learning in this setup but is agnostic to the physical system and energy function used, we refer to two recent works (Ernoult et al., ; Scellier et al., 2022).
@@ 3 Background
Let us consider a neural network with $L$ layers, and let us denote $\mathbf{W}^{l}$ and $\mathbf{x}^{l}_{t}$ the weight parameters and the neural activities of layer $l$, respectively. Note that, differently from standard models trained with BP, the neural activities are variables of the model, optimized over multiple time steps $t$. This optimization is performed with the goal of allowing the activities of every layer to predict those of the layer below. Together with the neural activities, the two other quantities usually related to single neurons are the prediction $\bf{\mu}^{l}_{t}=\mathbf{W}^{l}f\left(\mathbf{x}^{l-1}_{t}\right)$, given by the layer-wise operation through an activation function, and the prediction error, defined as the deviation of the actual activity from the prediction, that is, $\bm{\varepsilon}^{l}_{t}=\mathbf{x}^{l}_{t}-\mathbf{\mu}^{l}_{t}$. A fourth quantity, usually overlooked in machine learning applications but of vital importance in neuroscience, is the precision, or covariance $\Sigma^{l}_{t}$ of a specific neuron  [^1] In predictive coding, and more generally in statistics, the precision matrix is the matrix inverse of the covariance matrix. In this work, we follow the standard convention and divide the prediction error by a factor of $\Sigma$, instead of multiplying it by a factor of $p$.. Differently from the standard literature, however, we consider the covariance to be time-dependent. Furthermore, instead of updating it to minimize an objective function as done in previous works Ofner et al. (2021), we will manually define the rule that governs its updates. The predictive coding energy is then the sum of the squared norms of the precision-weighted prediction errors of every layer over time:
@@
where we consider covariances to be layer-dependent: all the neurons of the same layer will have the same covariance. To this end, we use the same notation when the covariance $\Sigma_{t}^{l}$ is a scalar value, or a diagonal matrix whose entries are equal to such a scalar. For a detailed representation of a predictive coding network, we refer to Figure 1(a).
@@
Suppose that we are provided with a labeled data point $(\mathbf{o},\mathbf{y})$, where $\mathbf{o}$ denotes sensory inputs, and $\mathbf{y}\in\mathbb{R}^{o}$ is the label. Training is then performed via a form of bi-level optimization (Zucchet and Sacramento, 2022), divided into three phases. In the first phase, the neural activities of every neuron are initialized via a forward pass, that is, we set $\bf x^{l}_{0}=\mu^{l}_{0}$ for every layer, with $x^{0}_{0}=o$. In the second phase, that we call the inference phase, we fix the neural activities of the output layer to the label, that is, we set $\mathbf{x}_{L}=\mathbf{y}$, and we update the neural activities via gradient descent, to minimize the total energy of the model. The update rule is then the following:
@@
where $\alpha$ is the learning rate of the neural activities. This phase will continue until it reaches the fixed number of iterations $\mathbf{T}$ or achieves convergence. The third phase is the learning phase, where the neural activities $\mathbf{x}^{l}_{T}$ are fixed, and the weight parameters are updated to decrease the energy, weighted by a learning rate $\eta$, via the following equation:
@@
Nudging as Precision.
Instead of providing the original label $y$ to the model, it is common in the literature to slightly translate the output neurons of the system $\mathbf{x}^{L}_{0}$ in the direction of $y$. More precisely, it fixes $\mathbf{x}^{L}_{t}=\mu^{L}_{0}+\beta(\mathbf{y}-\mu^{L}_{0})$ for every time step $t$, where $\beta$ controls the supervision strength. The sign of $\beta$ determines supervision polarity: positive for standard nudging and negative for inverse supervision. Performing a stochastic sampling from $\{\beta,-\beta\}$ across training epochs and batches is called center nudging (Scellier et al., 2024). In practice, the term $\beta$ scales the prediction error at the output layer, meaning that it functions as the "gain" or "precision" with which the error signal influences the state of the model, effectively controlling the strength and direction of the information flow from the desired output back into the system, much like precision weighting alters error unit gain and balances signals in predictive coding. To this end, it is possible to describe the same phenomenon using standard predictive coding and our dynamical precision weighting by simply setting $(\Sigma^{L}_{t})^{-1}=\beta$ for every $t$.

### v3

@@ Abstract
Predictive coding networks are neural models that perform inference through an iterative energy minimization process, whose operations are local in space and time. While effective in shallow architectures, they suffer significant performance degradation beyond five to seven layers. In this work, we show that this degradation is caused by exponentially imbalanced errors between layers during weight updates, and by predictions from the previous layers not being effective in guiding updates in deeper layers. Furthermore, when training models with skip connections, the energy propagated by the residuals reaches higher layers faster than that propagated by the main pathway, affecting test accuracy. We address the first issue by introducing a novel precision-weighted optimization of latent variables that balances error distributions during the relaxation phase, the second issue by proposing a novel weight update mechanism that reduces error accumulation in deeper layers, and the third one by using auxiliary neurons that slow down the propagation of the energy in the residual connections. Empirically, our methods achieve performance comparable to backpropagation on deep models such as ResNets, opening new possibilities for predictive coding in complex tasks.
@@ 1 Introduction
Training deep learning models is extremely expensive in terms of energy consumption. To address this problem, a recent direction of research is studying the use of alternative accelerators that leverage the properties of physical systems to perform computations (Wright et al., 2022).However, transitioning to new hardware without altering the main algorithm — error backpropagation (Rumelhart et al., 1986) — has proven to be challenging due to two central issues: the requirement for sequential forward and backward passes and the need to analytically compute gradients of a global cost function. These issues do not arise when using learning algorithms that rely on computations that are local in space and time (i.e., all updates rely on locally available information only) (Bengio, 2014; Hinton, 2022). A popular example is equilibrium propagation, a framework that allows for the learning of the parameters of a neural network by simulating a physical system brought to an equilibrium (Scellier & Bengio, 2017). This physical system is usually defined via an energy function that describes the state of a neural network in terms of its weights and neurons, with different functions describing different systems (Krotov & Hopfield, 2016).
@@
In recent years, researchers have devoted significant effort to scaling up the deployment of energy-based models. Two recent works have benchmarked multiple variants of common learning algorithms using the Hopfield energy function (Scellier et al., 2024) and the predictive coding energy (Pinchetti et al., 2024), showing that this class of models can match the performance of standard deep learning when training relatively shallow networks of up to five or seven layers. However, this success does not extend to deeper architectures, where performance degrades substantially. Notably, the degradation is even more severe in predictive coding implementations of residual networks, which perform worse than equally deep models without skip connections. Since the major advances of modern deep learning rely on very deep architectures with residual connections, understanding and overcoming these limitations is a critical step toward scaling predictive coding networks to the regimes where deep learning has been most successful.
@@
To understand the poor scalability of predictive coding networks, it is instructive to examine how energy propagates across depth. It has recently been shown that in a three-layer model, the energy concentrated in a layer can be up to an order of magnitude larger than the energy concentrated in the layer before (Pinchetti et al., 2024). While such shallow models can still achieve good test accuracies, we conjecture that this ’energy imbalance’ becomes a critical bottleneck in deeper architectures, leading to performance degradation in a way that is conceptually similar to the vanishing gradient problem (Hochreiter, 1998). More precisely, this imbalance prevents the effective propagation of energy — and crucially, the associated error information — from the output layer back to the early layers, creating two problems: first, it prevents the model from fully leveraging its depth, as early layers receive insufficient error signals for effective training; second, latent states may diverge substantially from their forward pass values due to the excessive energy in later layers.
Addressing this energy imbalance requires mechanisms that can adaptively modulate the relative influence of different layers as errors propagate through the network, a challenge similar to what biological neural systems solve through precision-based regulation. In predictive processing theories, precision refers to the estimated reliability of a prediction error (formally, its inverse variance), and it is thought to be dynamically regulated to balance bottom-up and top-down signals across cortical hierarchies (Feldman & Friston, 2010; Bastos et al., 2012). Despite its central role in biological inference, most machine learning formulations of predictive coding set precision to 1 for simplicity (Whittington & Bogacz, 2017), thereby overlooking its potential as a powerful mechanism to stabilize learning dynamics. The first contribution of this work is to propose leveraging precision weighting to regularize energy propagation in predictive coding networks.
We begin by analyzing the energy propagation of deep convolutional architectures for both predictive coding (PC) and incremental PC (iPC) — a recently introduced variant of PC that updates weights and neurons simultaneously Salvatori et al. (2024). Building on the insights from this analysis, we propose time-dependent precisions that address the identified issues. Our results show that this substantially improves test accuracy, supporting our conjecture of a causal link between energy propagation and empirical performance. More broadly, we propose two algorithmic improvements (spiking precision and a novel weight-update mechanism) and two structural improvements (PC-tailored batch normalization and auxiliary neurons for skip connections) that enable PC to achieve competitive performance with backprop on image classification benchmarks, including VGG models with up to $15$ layers and ResNets1818 on Tiny ImageNet. Our contributions can be detailed as follows:
  • We show that in models trained with PC, the energy is orders of magnitude larger in layers closer to the output, supporting the hypothesis that information fails to propagate effectively to the earlier layers. This phenomenon is less pronounced in iPC, where continuous weight updates rapidly reduce the excess energy. However, these updates result in even lower test accuracy.
  • To mitigate this imbalance, we propose dynamical precision-weightings that depend on both time and layer depth. The most effective variant, which we call spiking precision, applies very large precisions as soon as the energy reaches a given layer, thereby boosting it forward. Experiments show that this method regulates the energy imbalance and improves test accuracy in deep PC models. In the case of iPC, spiking precisions alone are already sufficient to achieve performance comparable to backpropagation in deep networks.
  • To further improve the performance of standard PC models, we introduce a novel weight-update mechanism that modifies how parameters are updated. This method combines predictions computed at initialization (hence adding a degree of implausibility, as they have to be stored in memory) with the neural activities at convergence, resulting in more effective updates. With this approach, PC attains performance on par with backpropagation and iPC on deep convolutional models. In addition, we propose a variant of batch normalization (Ioffe & Szegedy, 2015) tailored for PC, which further enhances performance.
  • While effective for VGG-like models, when training PC-based ResNets He et al. (2016), we still observe a significant drop in performance. We conjecture that this is caused by the energy propagated by the residuals, which reach higher layers faster than that propagated by the main pathway, disrupting learning dynamics. We show that this can be addressed by adding extra families of neurons inside the skip connections, that have the sole goal of slowing down the feedback signal of the skip connections so that it reaches the higher layers at the same time as the main one. The results show that such auxiliary neurons allow models trained with PC and iPC to reach performance comparable to these of backprop on ResNet18.
@@ 2 Related Works
EP is a learning algorithm for supervised learning that is largely inspired by contrastive learning on continuous Hopfield networks (Movellan, 1991). Here, neural activities are updated in two phases: In the first, to minimize an energy function defined on the parameters of the neural network; in the second, to minimize the same energy with the addition of a loss function defined on the labels (Scellier & Bengio, 2017). Interestingly, these two phases allow us to approximate the gradient of the loss function up to arbitrary levels of accuracy using finite difference coefficients (Zucchet & Sacramento, 2022). The consequence is that EP can be seen as a technique that allows the minimization of loss functions using arbitrary physical systems that can be brought to an equilibrium, and it has hence been studied in a large number of domains (Scellier, 2024; Kendall et al., 2020). The state of the art is that EP models are able to match the performance of BPTT (BP Through Time) on models with $5$ and $7$ hidden layers (Scellier et al., 2024), with the exception of hybrid models, which manage to achieve a good performance on models with $15$ layers by alternating blocks of layers trained with BP and blocks trained with EP (Nest & Ernoult, 2024).
@@
The formulation of PC that we use here was developed to model hierarchical information processing in the brain (Rao & Ballard, 1999; Friston, 2005). Intuitively, this theory states that neurons and synapses at one level of the hierarchy are updated to better predict the activities of the neurons of the layers below, and minimize the prediction error. Interestingly, the same algorithm can be used as a training algorithm for deep neural networks (Whittington & Bogacz, 2017), where several similarities with backpropagation were observed (Song et al., 2020; Salvatori et al., 2022). To this end, it has been used in a large number of machine learning tasks, from image generation and classification to natural language processing and associative memory (Sennesh et al., 2024; Salvatori et al., 2023; Pinchetti et al., 2022; Ororbia & Kifer, 2020; Salvatori et al., 2021). Again, the state of the art has been reached by training convolutional models with $5$ hidden layers, with performance starting to get worse as soon as we use models with $7$ layers (Pinchetti et al., 2024). To this end, a recent interesting work has proposed $\mu PC$, a theoretical framework that allows the training of very deep feedforward models Innocenti et al. (2025). However, this work does not tackle non-feedforward layers, and does not test on datasets larger than MNIST. The connection between PC and EP is well explained by the concept of bi-level optimization (Zucchet & Sacramento, 2022), where the neural activities used for learning are the equilibrium state of a physical system.
@@ 3 Background
Let us consider a network with $L$ layers, and let us denote $\mathbf{W}^{l}$ and $\mathbf{x}^{l}_{t}$ the weight parameters and the neural activities of layer $l$, respectively. Note that, differently from standard models, the neural activities are variables of the model, optimized over time steps $t$. This optimization is performed with the goal of allowing the activities of every layer to predict those of the layer below. Together with the neural activities, the two other quantities related to single neurons are the prediction $\bf{\mu}^{l}_{t}=\mathbf{W}^{l}f\left(\mathbf{x}^{l-1}_{t}\right)$, given by the layer-wise operation through an activation function, and the prediction error, defined as the deviation of the actual activity from the prediction, that is, $\bm{\varepsilon}^{l}_{t}=\mathbf{x}^{l}_{t}-\mathbf{\mu}^{l}_{t}$. A fourth quantity, usually overlooked in machine learning applications but of vital importance in neuroscience, is the precision, or covariance $\Sigma^{l}_{t}$ of a specific neuron  [^1] In predictive coding, and more generally in statistics, the precision matrix is the matrix inverse of the covariance matrix. In this work, we follow the standard convention and divide the prediction error by a factor of $\Sigma$, instead of multiplying it by a factor of $p$.. Differently from the standard literature, we consider the covariance to be time-dependent. Furthermore, instead of updating it to minimize an objective function as done in previous works (Ofner et al., 2021), we will manually define the rule that governs its updates. The predictive coding energy is then the sum of the squared norms of the precision-weighted prediction errors of every layer over time:
@@
where we consider covariances to be layer-dependent: all the neurons of the same layer will have the same covariance. To this end, we use the same notation when the covariance $\Sigma_{t}^{l}$ is a scalar, or a diagonal matrix whose entries are equal to such a scalar.
@@
Suppose that we are provided with a labeled data point $(\mathbf{o},\mathbf{y})$, where $\mathbf{o}$ denotes sensory inputs, and $\mathbf{y}\in\mathbb{R}^{o}$ is the label. Training is then performed via a form of bi-level optimization (Zucchet & Sacramento, 2022), divided into three phases. In the first phase, the neural activities of every neuron are initialized via a forward pass, that is, we set $\bf x^{l}_{0}=\mu^{l}_{0}$ for every layer, with $x^{0}_{0}=o$. In the second phase, that we call the inference phase, we fix the neural activities of the output layer to the label, that is, we set $\mathbf{x}_{L}=\mathbf{y}$, and we update the neural activities via gradient descent, to minimize the total energy of the model. The update rule is then the following:
@@
where $\alpha$ is the learning rate of the neural activities. This phase will continue until it reaches the fixed number of iterations $\mathbf{T}$ or convergence. The third phase is the learning phase, where the neural activities $\mathbf{x}^{l}_{T}$ are fixed, and the weight parameters are updated to decrease the energy, weighted by a learning rate $\eta$, via the following equation:
@@
Incremental PC (iPC).
An alternative to the bi-level optimization described above is iPC Salvatori et al. (2024), that differs from the standard implementation as also the weight parameters are updated at every time step $t$ until convergence, according to Eq.3. From the variational inference perspective, this schema introduces a form of incremental expectation-maximization Neal & Hinton (1998), and has been argued to be more biologically plausible than standard PC, as it avoids the need for a control signal that halts the updates of the neural activities and triggers the weight updates. In practice, however, this method has been shown to perform even worse than PC when it is used to train deep models such as VGG7 and VGG9 on CIFAR10 Pinchetti et al. (2022). We will show that it is this fully automatic training method that will benefit the most from hard-coded precisions, and reach the best performance on models such as ResNet18.
Nudging.
Instead of providing the original label $y$ to the model, it is common in the literature to slightly translate the output neurons of the system $\mathbf{x}^{L}_{0}$ in the direction of $y$. More precisely, it fixes $\mathbf{x}^{L}_{t}=\mu^{L}_{0}+\beta(\mathbf{y}-\mu^{L}_{0})$ for every time step $t$, where $\beta$ controls the supervision strength. The sign of $\beta$ determines supervision polarity: positive for standard nudging and negative for inverse supervision. Performing a stochastic sampling from $\{\beta,-\beta\}$ across training epochs and batches is called center nudging (Scellier et al., 2024). In practice, PC with centered nudging has been shown to be the best performing method on deep models such as VGG7 Pinchetti et al. (2024).